# PPFPL: Cross-silo Privacy-preserving Federated Prototype Learning Against Data Poisoning Attacks on Non-IID Data

### v4

@@ 6       $b_{n+1}=b_{n}^{2}(\frac{b_{n}-3}{4})$;
where $\textit{Round}(p\times h^{k}_{m,t},6)$ denotes that $p\times h^{k}_{m,t}$ is rounded to the 6-th decimal place. This is because CKKS decrypts the ciphertext with an error in the range of $10^{-7}$ [50]. Moreover, $\textit{min}^{k}_{t}$ denotes the smallest value for class $k$ in set $\{\textit{Round}(p\times h^{k}_{m,t},6)\}_{m\in\mathcal{S}}^{|\mathcal{S}|}$. The $\textit{Round}(p\times h^{k}_{m,t},6)=\textit{min}^{k}_{t}$ means that the local prototype $\textit{{c}}^{k}_{m,t}$ satisfies $\textit{sim}^{k}_{m,t}<\chi$, thus its aggregation weight $j^{k}_{m,t}$ is 0. Additionally, Verifier calculates $\textit{Sum}^{k}_{t}=\sum_{m\in\mathcal{S}}j^{k}_{m,t}$ and encrypts $j^{k}_{m,t}$ and $\textit{{V}}\odot\widetilde{\textit{{c}}}^{k}_{m,t}$ using the client’s public key $Pkx$ to get $\llbracket j^{k}_{m,t}\rrbracket_{Pkx}$ and $\llbracket\textit{{V}}\odot\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkx}$, and sends them to Aggregator.
The Aggregator computes $\frac{1}{\textit{{V}}}\odot\llbracket\textit{{V}}\odot\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkx}$ to get $\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkx}$, and aggregates the encrypted local prototypes to get the encrypted global prototype $\llbracket\textit{{C}}^{k}_{t+1}\rrbracket_{Pkx}$ by the following formula,
@@
VI Analysis
In the section, we provide both convergence analysis and privacy analysis for PPFPL. Specifically, we make the following assumptions similar to existing general frameworks [22][51] for loss function (3).
@@ VI Analysis
where $\mathcal{L}_{m,t}^{(e_{1})}$ denotes loss function at the $(tE+e_{1})$-th local iteration in client $m$. This implies the following quadratic bound,
$\mathcal{L}_{m,t}^{(e_{1})}$ $-\mathcal{L}_{m,t}^{(e_{2})}$
$\leq\left\langle\nabla\mathcal{L}_{m,t}^{(e_{2})},\left(\textit{{w}}_{m,t}^{(e_{1})}-\textit{{w}}_{m,t}^{(e_{2})}\right)\right\rangle+\frac{L_{1}}{2}\left\|\textit{{w}}_{m,t}^{(e_{1})}-\textit{{w}}_{m,t}^{(e_{2})}\right\|_{2}^{2}.$
@@
and its variance is bounded by $\sigma^{2}$:
$\mathbb{E}[\|\textit{{g}}_{m,t}^{(e)}-\nabla\mathcal{L}(\textit{{w}}_{m,t}^{(e)})\|_{2}^{2}]\leq\sigma^{2}.$
Based on the above assumptions, we have the following theorem and corollaries. Notably, we add “$\frac{1}{2}$” into the local iteration, denoted as $\{\frac{1}{2},1,\cdots,E\}$ in our analysis. For example, $tE$ denotes the time step before local prototype aggregation, and $tE+\frac{1}{2}$ denotes the time step between local prototype aggregation and the first local iteration of the $t$-th round.
@@
In PPFPL, regardless of the percentage of malicious clients, for the $t$-th communication round, the variation of loss function for each benign client can be bounded as,
$\mathbb{E}$ $\left[\mathcal{L}_{m,t+1}^{\frac{1}{2}}\right]-\mathcal{L}_{m,t}^{E}\leq G(\lambda,\eta,E),$
where
$G(\lambda,\eta,E)=-\left(\eta-\frac{\eta^{2}L_{1}}{2}\right)\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}+\frac{E\eta^{2}L_{1}}{2}\sigma^{2}+2\lambda.$
@@
$\mathcal{L}_{m,t}^{(1)}$ $\leq\mathcal{L}_{m,t}^{(\frac{1}{2})}+\left\langle\nabla\mathcal{L}_{m,t}^{(\frac{1}{2})},\left(\textit{{w}}_{m,t}^{(1)}-\textit{{w}}_{m,t}^{(\frac{1}{2})}\right)\right\rangle+\frac{L_{1}}{2}\left\|\textit{{w}}_{m,t}^{(1)}-\textit{{w}}_{m,t}^{(\frac{1}{2})}\right\|_{2}^{2}$ (10)
@@
$\mathbb{E}[\mathcal{L}_{m,t}^{(1)}]\leq\mathcal{L}_{m,t}^{(\frac{1}{2})}-\eta\mathbb{E}\left[\left\langle\nabla\mathcal{L}_{m,t}^{(\frac{1}{2})},\textit{{g}}_{m,t}^{(\frac{1}{2})}\right\rangle\right]+\frac{\eta^{2}L_{1}}{2}\mathbb{E}\left[\left\|\textit{{g}}_{m,t}^{(\frac{1}{2})}\right\|_{2}^{2}\right]$
@@
[…]
$=\mathcal{L}_{m,t}^{(\frac{1}{2})}-(\eta-\frac{\eta^{2}L_{1}}{2})\|\nabla\mathcal{L}_{m,t}^{(\frac{1}{2})}\|_{2}^{2}+\frac{\eta^{2}L_{1}}{2}Var(\textit{{g}}_{m,t}^{(\frac{1}{2})})$
[…]
where (b), (d) and (e) follow from Assumption 2, (c) follows from $Var(x)=\mathbb{E}[x^{2}]-(\mathbb{E}[x])^{2}$. Then, during the local computation step, the loss function is updated $E$ times, and it can be bounded as:
$\mathbb{E}[\mathcal{L}_{m,t}^{(1)}]\leq\mathcal{L}_{m,t}^{(\frac{1}{2})}-(\eta-\frac{\eta^{2}L_{1}}{2})\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}+\frac{E\eta^{2}L_{1}}{2}\sigma^{2}.\$ (11)
@@
$\mathcal{L}_{m,t+1}^{\frac{1}{2}}=\mathcal{L}_{m,t}^{E}+\mathcal{L}_{m,t+1}^{\frac{1}{2}}-\mathcal{L}_{m,t}^{E}$ (12)
$\stackrel{{\scriptstyle}}{{=}}\mathcal{L}_{m,t}^{E}+\lambda\mathcal{L_{R}}(\textit{{c}}^{k}_{m,t+1},\textit{{C}}^{k}_{t+2})-\lambda\mathcal{L_{R}}(\textit{{c}}^{k}_{m,t+1},\textit{{C}}^{k}_{t+1})$
$\overset{}{=}\mathcal{L}_{m,t}^{E}-\frac{\lambda}{|\mathcal{I}|}\sum_{k\in\mathcal{I}}\textit{sim}(\textit{{c}}_{m,t+1}^{k},\textit{{C}}^{k}_{t+2})+\frac{\lambda}{|\mathcal{I}|}\sum_{k\in\mathcal{I}}\textit{sim}(\textit{{c}}_{m,t+1}^{k},\textit{{C}}^{k}_{t+1})$
[…]
where (f) follows from $-1\leq\textit{sim}(\cdot,\cdot)\leq 1$. Taking expectation on both sides of formula (12), we can get
$\mathbb{E}\left[\mathcal{L}_{m,t+1}^{\frac{1}{2}}\right]$ $\leq\mathcal{L}_{m,t}^{E}+2\lambda.$ (13)
Thus, during the $t$-th communication round, according to the formula (11) and formula (13), the variation of loss function for each benign client can be bounded as,
$\mathbb{E}$ $\left[\mathcal{L}_{m,t+1}^{\frac{1}{2}}\right]-\mathcal{L}_{m,t}^{E}\leq G(\lambda,\eta,E),$
where
$G(\lambda,\eta,E)=-\left(\eta-\frac{\eta^{2}L_{1}}{2}\right)\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}+\frac{E\eta^{2}L_{1}}{2}\sigma^{2}+2\lambda.$
Thus, Theorem 1 is proved. ∎
@@
Given any fixed $\lambda$ and $E$, the $G(\eta)$ is a convex function with respect to $\eta$.
@@
To prove that $G(\eta)$ is a convex function for $\eta$, we need to prove that the second order derivative of $G(\eta)$ with respect to $\eta$ is always nonnegative. Thus, we have
$\frac{dG(\eta)}{d\eta}=-\left(1-L_{1}\eta\right)\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}+L_{1}E\eta\sigma^{2},$
@@
$\frac{d^{2}G(\eta)}{d\eta^{2}}=L_{1}\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}+L_{1}E\sigma^{2}.$
@@
When $G^{\prime}(\eta)$ equals 0, then $G(\eta)$ obtains a extremum value. Let $(\frac{dG(\eta)}{d\eta}|\eta=\eta^{\star})=0,$ so we get
@@
Thus, $G(\eta)$ exists a extremum value when $\eta=\eta^{\star}=\frac{\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}}{L_{1}E\sigma^{2}+L_{1}\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}}.$ Because corollary 1 proves the second order derivative of $G(\eta)$ with respect to $\eta$ is always nonnegative, the extremum value is the minimal value. Therefore, we can understand that the variation of loss function exists a minimum bound when $\eta=\eta^{\star}$. ∎
@@
To guarantee that the local loss function decreases after each communication, we need to make sure that $G(\lambda)\textless 0$, so we have
$-\left(\eta-\frac{L_{1}\eta^{2}}{2}\right)\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}+\frac{E\eta^{2}L_{1}}{2}\sigma^{2}+2\lambda\textless 0.$
After simplification, we get
$\lambda\textless\left(\frac{1}{2}\eta-\frac{L_{1}\eta^{2}}{4}\right)\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}-\frac{E\eta^{2}L_{1}}{4}\sigma^{2}.$
@@
During the secure aggregation, the two servers can obtain plaintext information $\|\textit{{C}}^{\prime k}_{t+1}\|$ and $\textit{Sum}^{k}_{t}$, where $\|\textit{{C}}^{\prime k}_{t+1}\|$ denotes the norm of trusted update, $\textit{Sum}^{k}_{t}$ denotes the sum of aggregation weight of class $k$. For non-colluding Aggregator and Verifier, they cannot get any sensitive information from these plaintext information. In addition, for colluding malicious clients, when there are $(|\mathcal{S}|-1)$ malicious clients in PPFPL, they can infer the local prototype about benign client from the encrypted global prototype. However, the real scenario does not exist when there are $(|\mathcal{S}|-1)$ malicious clients. Since benign clients receive only global prototypes distributed by Aggregator, they cannot deduce information about others. Therefore, neither third-party entities nor malicious clients can derive sensitive information about benign clients. ∎
VII Experiments

### v5

@@ 6  $b_{n+1}=b_{n}^{2}(\frac{b_{n}-3}{4})$;
where $\textit{Round}(p\times h^{k}_{m,t},6)$ denotes that $p\times h^{k}_{m,t}$ is rounded to the 6-th decimal place. This is because CKKS decrypts the ciphertext with an error in the range of $10^{-7}$ [53]. Moreover, $\textit{min}^{k}_{t}$ denotes the smallest value for class $k$ in set $\{\textit{Round}(p\times h^{k}_{m,t},6)\}_{m\in\mathcal{S}}^{|\mathcal{S}|}$ . The condition $\textit{Round}(p\times h^{k}_{m,t},6)=\textit{min}^{k}_{t}$ means that the local prototype $\textit{{c}}^{k}_{m,t}$ satisfies $\textit{sim}^{k}_{m,t}<\chi$, thus its aggregation weight $j^{k}_{m,t}$ is 0. Additionally, Verifier calculates $\textit{Sum}^{k}_{t}=\sum_{m\in\mathcal{S}}j^{k}_{m,t}$ and encrypts $j^{k}_{m,t}$ and $\textit{{V}}\odot\widetilde{\textit{{c}}}^{k}_{m,t}$ using the client’s public key $Pkx$ to get $\llbracket j^{k}_{m,t}\rrbracket_{Pkx}$ and $\llbracket\textit{{V}}\odot\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkx}$ , and sends them to Aggregator.
The Aggregator computes $\frac{1}{\textit{{V}}}\odot\llbracket\textit{{V}}\odot\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkx}$ to get $\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkx}$ , and aggregates the encrypted local prototypes to get the encrypted global prototype $\llbracket\textit{{C}}^{k}_{t+1}\rrbracket_{Pkx}$ by the following formula:
@@
6 Analysis
In the section, we provide both convergence analysis and privacy analysis for PPFPL. Specifically, we make the following assumptions similar to existing general frameworks [23][54] for the local optimization function (3).
@@ 6 Analysis
where $\mathcal{L}_{m,t}^{(e_{1})}$ denotes loss function at the $(tE+e_{1})$-th local iteration in client $m$. This implies the quadratic bound:
$\begin{aligned} \mathcal{L}_{m,t}^{(e_{1})}&-\mathcal{L}_{m,t}^{(e_{2})}\leq\left\langle\nabla\mathcal{L}_{m,t}^{(e_{2})},\left(\textit{{w}}_{m,t}^{(e_{1})}-\textit{{w}}_{m,t}^{(e_{2})}\right)\right\rangle+\frac{L_{1}}{2}\left\|\textit{{w}}_{m,t}^{(e_{1})}-\textit{{w}}_{m,t}^{(e_{2})}\right\|_{2}^{2}.\end{aligned}$
@@
and its variance is bounded by $\sigma^{2}$: $\mathbb{E}[\|\textit{{g}}_{m,t}^{(e)}-\nabla\mathcal{L}(\textit{{w}}_{m,t}^{(e)})\|_{2}^{2}]\leq\sigma^{2}$ .
Based on the above assumptions, we formulate the following theorem and corollaries. Notably, we add “ $\frac{1}{2}$ ” into the local iteration, denoted as $\{\frac{1}{2},1,\cdots,E\}$ in our analysis. For example, $tE$ denotes the time step before local prototype aggregation, and $tE+\frac{1}{2}$ denotes the time step between local prototype aggregation and the first local iteration in the $t$-th round.
@@
In PPFPL, under any proportion of malicious clients, for the $t$-th communication round, the variation of loss function for each benign client can be bounded as:
$\begin{aligned} \mathbb{E}&\left[\mathcal{L}_{m,t+1}^{\frac{1}{2}}\right]-\mathcal{L}_{m,t}^{E}\leq G(\lambda,\eta,E),\end{aligned}$
where $\begin{aligned} G(\lambda,\eta,E)=-\left(\eta-\frac{\eta^{2}L_{1}}{2}\right)\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}+\frac{E\eta^{2}L_{1}}{2}\sigma^{2}+2\lambda.\end{aligned}$
We can observe that each variable in this upper bound is independent of the malicious clients, indicating that the impact of malicious clients is constrained. This constraint arises from incorporating the cosine similarity in the auxiliary term of our local optimization function, which restricts the update direction of the local model.
@@
where (b), (d) and (e) are derived from Assumption 2, (c) follows from $Var(x)=\mathbb{E}[x^{2}]-(\mathbb{E}[x])^{2}$ . During the local computation step, the loss function is updated $E$ times, and it can be bounded as:
$\mathbb{E}[\mathcal{L}_{m,t}^{(1)}]\leq\mathcal{L}_{m,t}^{(\frac{1}{2})}-(\eta-\frac{\eta^{2}L_{1}}{2})\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}+\frac{E\eta^{2}L_{1}}{2}\sigma^{2}.\\$
@@
$\begin{aligned} &\mathcal{L}_{m,t+1}^{\frac{1}{2}}=\mathcal{L}_{m,t}^{E}+\mathcal{L}_{m,t+1}^{\frac{1}{2}}-\mathcal{L}_{m,t}^{E}\\ &\stackrel{{\scriptstyle}}{{=}}\mathcal{L}_{m,t}^{E}+\lambda\mathcal{L_{R}}(\textit{{c}}^{k}_{m,t+1},\textit{{C}}^{k}_{t+2})-\lambda\mathcal{L_{R}}(\textit{{c}}^{k}_{m,t+1},\textit{{C}}^{k}_{t+1})\\ &\overset{}{=}\mathcal{L}_{m,t}^{E}-\frac{\lambda}{|\mathcal{I}|}\sum_{k\in\mathcal{I}}\textit{sim}(\textit{{c}}_{m,t+1}^{k},\textit{{C}}^{k}_{t+2})+\frac{\lambda}{|\mathcal{I}|}\sum_{k\in\mathcal{I}}\textit{sim}(\textit{{c}}_{m,t+1}^{k},\textit{{C}}^{k}_{t+1})\\ &\overset{\text{(f)}}{\leq}\mathcal{L}_{m,t}^{E}+2\lambda,\end{aligned}$
where (f) follows from $1\leq\textit{sim}(\cdot,\cdot)\leq 1$ . Notably, although different proportions of malicious clients can influence the direction of $\textit{{C}}^{k}_{t+1}$, the contribution of malicious prototypes is inherently bounded by the cosine similarity $\textit{sim}(\cdot,\cdot)$. In other words, regardless of the proportion of the malicious clients, PPFPL suppresses the impact of malicious clients, thereby maintaining the robustness of the aggregated results.
Taking expectation on both sides of formula (12), we can get:
$\begin{aligned} \mathbb{E}\left[\mathcal{L}_{m,t+1}^{\frac{1}{2}}\right]&\leq\mathcal{L}_{m,t}^{E}+2\lambda.\end{aligned}$
Thus, during the $t$-th communication round, according to the formula (11) and formula (13), the variation of loss function for each benign client can be bounded as:
$\begin{aligned} \mathbb{E}&\left[\mathcal{L}_{m,t+1}^{\frac{1}{2}}\right]-\mathcal{L}_{m,t}^{E}\leq G(\lambda,\eta,E),\end{aligned}$
where $G(\lambda,\eta,E)=-\left(\eta-\frac{\eta^{2}L_{1}}{2}\right)\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}+\frac{E\eta^{2}L_{1}}{2}\sigma^{2}+2\lambda$ . Thus, Theorem 1 is proved. ∎
@@
Given any fixed $\lambda$ and $E$, the $G(\eta)$ is convex with respect to $\eta$.
@@
To prove that $G(\eta)$ is convex for $\eta$, we need to prove that the second order derivative of $G(\eta)$ with respect to $\eta$ is always nonnegative. Thus, we have
$\begin{aligned} \frac{dG(\eta)}{d\eta}=-\left(1-L_{1}\eta\right)\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}+L_{1}E\eta\sigma^{2},\end{aligned}$
@@
$\begin{aligned} \frac{d^{2}G(\eta)}{d\eta^{2}}=L_{1}\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}+L_{1}E\sigma^{2}.\end{aligned}$
@@
When $G^{\prime}(\eta)$ equals 0, then $G(\eta)$ obtains a extremum value. Let $(\frac{dG(\eta)}{d\eta}|\eta=\eta^{\star})=0$ , we can get:
@@
Thus, $G(\eta)$ exists a extremum value when $\eta=\eta^{\star}=\frac{\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}}{L_{1}E\sigma^{2}+L_{1}\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}}$ . Because corollary 1 proves the second order derivative of $G(\eta)$ with respect to $\eta$ is always nonnegative, the extremum value is the minimal value. Therefore, we can understand that it exists a minimum bound when $\eta=\eta^{\star}$. ∎
@@
To guarantee that the local loss function decreases after each communication, it is necessary to make sure that $G(\lambda)\textless 0$. We have:
$\begin{aligned} -\left(\eta-\frac{L_{1}\eta^{2}}{2}\right)\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}+\frac{E\eta^{2}L_{1}}{2}\sigma^{2}+2\lambda\textless 0.\end{aligned}$
After simplification, we can get:
$\begin{aligned} \lambda\textless\left(\frac{1}{2}\eta-\frac{L_{1}\eta^{2}}{4}\right)\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}-\frac{E\eta^{2}L_{1}}{4}\sigma^{2}.\end{aligned}$
Theorem 2.
Let the proportions of benign and malicious clients be $(1-\kappa)$ and $\kappa$, respectively, with $\kappa\in[0,1]$. In the $t$-th round, let $\textit{{u}}_{ben}^{k}$ denote the average prototype of benign clients for class $k$, and $\textit{{u}}_{mal}^{k}$ denote the average prototype of malicious clients for class $k$. Then, the aggregated global prototype for class $k$ is given by:
$\textit{{C}}^{k}_{t+1}=(1-\kappa)\textit{{u}}_{ben}^{k}+\kappa\textit{{u}}_{mal}^{k}$
and its deviation in magnitude from the benign prototypes satisfies:
$\|\textit{{C}}^{k}_{t+1}-\textit{{u}}_{ben}^{k}\|_{2}=\kappa\|\textit{{u}}_{mal}^{k}-\textit{{u}}_{ben}^{k}\|_{2}\leq 2\kappa.$
From the above formula, we can observe that the deviation of the global prototype from the benign prototype increases linearly with the malicious proportion $\kappa$, thus ensuring that the influence of malicious clients always remains within a controllable range.
Proof.
By definition of the malicious client proportion $\kappa$, the global prototype can be expressed as:
[…]
Subtracting $\textit{{u}}_{ben}^{k}$ from both sides yields:
$\begin{aligned} \textit{{C}}^{k}_{t+1}-\textit{{u}}_{ben}^{k}=\kappa(\textit{{u}}_{mal}^{k}-\textit{{u}}_{ben}^{k}).\end{aligned}$
Taking the $\ell_{2}$-norm on both sides, we have:
$\begin{aligned} \|\textit{{C}}^{k}_{t+1}-\textit{{u}}_{ben}^{k}\|_{2}=\kappa\|(\textit{{u}}_{mal}^{k}-\textit{{u}}_{ben}^{k})\|_{2}.\end{aligned}$
Since $\textit{{u}}_{mal}^{k}$ and $\textit{{u}}_{ben}^{k}$ are unit vectors, the triangle inequality gives:
$\begin{aligned} \|\textit{{u}}_{mal}^{k}-\textit{{u}}_{ben}^{k}\|_{2}\leq\|\textit{{u}}_{mal}^{k}\|+\|\textit{{u}}_{ben}^{k}\|=2.\end{aligned}$
Consequently, we obtain the relation $\|\textit{{C}}^{k}_{t+1}-\textit{{u}}_{ben}^{k}\|_{2}\leq 2\kappa.\$This completes the proof of Theorem 2. ∎
@@
During the secure aggregation, the two servers can obtain plaintext values $\|\textit{{C}}^{\prime k}_{t+1}\|$ and $\textit{Sum}^{k}_{t}$, where $\|\textit{{C}}^{\prime k}_{t+1}\|$ denotes the norm of trusted prototype, $\textit{Sum}^{k}_{t}$ denotes the sum of aggregation weight of class $k$. For non-colluding Aggregator and Verifier, they cannot get any sensitive information from these plaintext information. In addition, for colluding malicious clients, if $(|\mathcal{S}|-1)$ clients are compromised, they can theoretically infer the local prototype of the remaining benign from the encrypted global prototype. However, the real scenario does not exist when there are $(|\mathcal{S}|-1)$ malicious clients. Since benign clients receive only global prototypes distributed by Aggregator, they cannot deduce information about others. Therefore, neither third-party entities nor malicious clients can derive sensitive information about benign clients. ∎
7 Experiments